OpenStudy (anonymous):

A team of engineers must design a fuel tank in the shape of a cone (excluding the base) is given by the formula S=Pi x square root of (r^2 + h^2). Find the radius of a cone with a height of 21 meters and a surface area of 155 meters squared. Round your answer to the nearest tenth if necessary

OpenStudy (irishboy123):

\[S = \pi (r^2 + h^2)\] i take it you want r for a given S, h combination, so you need to separate out r from the equation so that it reads \(r = \dots\) so you can start with \[ \dfrac{S}{\pi} = (r^2 + h^2)\] can you take if from there?